solve the equation: 3/x-5/=12
I am a home-schooled student and I am re-taking a class that I have forgotten everything in...

Answer by ankor@dixie-net.com(12692)

:
What does "3/x-5/=12" mean?
:
Got your message, use the uppercase back slash:
:
3|x-5| = 12;
:
Divide both sides by 3 and you have:
|x - 5| = 12/3
:
|x - 5| = 4
:
Remove the absolute signs and we have two equations with two solutions:
1st equation:
x - 5 = +4
x = 4 + 5; added 5 to both sides:
x = +9
:
2nd equation:
x - 5 = -4
x = -4 + 5; added 5 to both side:
x = +1
:
:
Check both solution in the original equation
x = +1
3|1 - 5| =
3|-4| =
3 * +4 = 12; everything inside the absolute comes out positive, right?
and
x = + 9
3|9 - 5| =
3|+4| =
3 * 4 = 12 also
:
Let me know if you receive this edited solution